# Large-scale moral machine experiment on large language models

### v1

@@ Results
To identify factors influencing the distance from human judgments, we first examined differences based on model types. This investigation was motivated by the general assumption that proprietary models demonstrate superior performance compared to open-source models [50], suggesting potential differences in their alignment with human judgments.
Models were classified into three groups: proprietary models, open-source models, and open-source models with parameters exceeding 10B (Fig 3). Proprietary models showed a median distance of 0.9 from human judgments, while open-source models as a whole demonstrated larger distances with a median of 1.2 (Wilcoxon test’s $p$-value $p=0.0081$).
However, to determine whether this difference was fundamentally attributable to the proprietary/open-source distinction or merely reflected differences in model sizes, we conducted additional analysis focusing on open-source models with parameters exceeding 10B. Results showed that large-scale open-source models exhibited a median distance of 0.9 from human judgments, with no significant difference from proprietary models (Wilcoxon test’s $p=0.93$).
These findings suggest that differences in judgment distance from humans depend more strongly on model size rather than the proprietary/open-source distinction.
@@
We conducted a detailed analysis of the relationship between model size and distance from human judgments, focusing on open-source models (Fig 4). Results revealed a significant negative correlation between model size and distance from human judgments (Spearman’s rank correlation coefficient $\rho=-0.50$, $p=0.018$), indicating that larger models tend to align more closely with human judgments.
Particularly in the Llama family, while 1-3B scale models (Llama 3.2 1B, 3B) showed distances of 1.2-1.5, 70B scale models (Llama 3 70B, 3.1 70B) demonstrated notably smaller distances of 0.7. Similar trends were observed in the Gemma family, where 2-9B scale models from Gemma 1.1 onward showed relatively large distances of 1.2-1.9, while the Gemma 2 27B scale model demonstrated a smaller distance of 1.1.
However, as indicated by the moderate correlation coefficient, the relationship between model size and distance from human judgments is not straightforward. For instance, within the Llama family, while Llama 3.2 1B showed a distance of 1.2, Llama 3.2 3B demonstrated a distance of 1.5. Similarly, among 8B models, version 3 showed a distance of 0.8, while version 3.1 exhibited a distance of 1.6. In the Gemma family, while Gemma 2B showed the minimum distance, subsequent larger models demonstrated increased distances. These observations suggest that while model size is a significant factor influencing judgment distance from humans, other factors (such as model architecture and training methodology) may play equally important roles.
Note that proprietary models were excluded from this analysis due to limited information about their sizes. However, given that proprietary models are generally believed to have parameters exceeding several hundred billion and demonstrate small median distances of 0.9 from human judgments as previously mentioned, this correlation would likely persist even if proprietary models were included.
@@
While previous analyses suggested model size as a significant factor influencing distance from human judgments, our observations indicated that updates within the same model family do not necessarily reduce judgment distances. Therefore, we conducted a more detailed analysis of how model updates affect distances from human judgments.
The impact of updates on open-source LLMs can be observed in detail from Fig 4. For instance, examining 2B Gemma models revealed increasing distances from human judgments through version progression from Gemma 2B to 1.1 2B and 2 2B. Similar trends were observed in 7–9B Gemma models, suggesting that updates do not necessarily lead to closer alignment with human judgment patterns.
The Llama family also demonstrated non-uniform effects of updates. Among 7–8B models, while the update from Llama 2 to 3 decreased distance from 1.1 to 0.8, showing closer alignment with human judgments, the subsequent update to Llama 3.1 increased distance to 1.6. Conversely, larger 70B models maintained a consistent distance of 0.7 through the update from Llama 3 to 3.1.
For proprietary models, we conducted temporal analysis of the relationship between release dates and distances from human judgments (Fig 5), focusing on GPT-3.5, GPT-4, GPT-4o, Gemini Pro, Gemini Flash, and Claude Sonnet families, which underwent multiple updates.
The GPT-3.5 and GPT-4 families maintained distances between 0.6–0.9 from March 2023 to April 2024. GPT-4 achieved its minimum distance of 0.6 with the January 2024 version but increased to 0.8 in the April version. GPT-3.5 showed its minimum distance of 0.7 in the June 2023 version, increased slightly to 0.9 in the November 2023 version, then returned to 0.8 in January 2024.
The newer GPT-4o family (from May 2024) showed slightly larger distances of 0.9–1.0 compared to traditional GPT-4. The Gemini family maintained relatively small distances of 0.9 through 1.0 Pro to 1.5 Pro Preview (April 2024), but later versions increased to 1.1. Gemini 1.5 Flash showed large initial distances of 1.2 (Pre, 001 versions) but decreased to 0.9 in version 002. The Claude Sonnet family’s distance decreased from 1.2 to 0.9 with the update from 3 to 3.5 (June 2024) but increased to 1.0 with the October 2024 update of 3.5. The Claude Haiku family’s distance increased from 0.9 to 1.1 with the update from 3 to 3.5.
These observations indicate that updates, whether in open-source or proprietary models, do not necessarily lead to closer alignment with human judgments. Furthermore, newer versions or supposedly higher-performing models do not consistently demonstrate closer alignment with human judgments.
@@
We conducted PCA using AMCE values across nine categories as feature vectors to comprehensively evaluate similarities in judgment tendencies among models (Fig 6). The first and second principal components explained 74.2% of total variance (46.1% and 28.1% respectively), enabling visualization of judgment similarities in this two-dimensional plane.
@@
Interestingly, these results aligned with our previous distance analysis in multiple aspects. Models showing smaller distances from human judgments tended to position closer to humans in this two-dimensional plane. Similar trends were observed regarding model size effects. For instance, within the Llama family, 70B models (Llama 3 70B, 3.1 70B) positioned near humans, while smaller 1-8B models distributed farther away.
@@ Discussion
In this study, we significantly expanded upon the four LLMs analyzed in our previous study [33] to comprehensively examine ethical judgments in autonomous driving across 51 different LLMs using the Moral Machine experimental framework (Figs 1 and 2). Proprietary models and open-source models exceeding 10B parameters demonstrated relatively small distances from human judgments (Figs 3 and 4), showing consistent alignment with human preferences in fundamental moral decisions, such as prioritizing humans over animals and favoring the preservation of more lives. However, model updates did not necessarily lead to closer alignment with human judgments, and variations in judgment tendencies were observed across updates (Fig 5). This extensive model comparison enabled a more comprehensive understanding of the evolution of moral judgments in LLMs.
@@
The incomplete alignment between LLM and human moral judgments is an expected outcome, considering that these models are not explicitly trained to mimic human moral judgments. However, the significant negative correlation observed between model size and distance from human judgments in open-source models (Fig 4) warrants attention. This finding aligns with scaling laws widely observed in language model behavior, suggesting that merely increasing model size might enable LLMs to naturally acquire human-like moral judgments. Nevertheless, this correlation does not necessarily imply causation, and other factors such as differences in model architecture, training data quality, and optimization methods must be considered. For instance, we observed cases within the same model family where size increases did not result in monotonic decreases in distance from human judgments.
@@
Differences in judgment tendencies across model families provide important implications for implementing autonomous driving systems. While the GPT-4 family demonstrated the closest alignment with human judgments (distances 0.6–0.8), its proprietary nature and black-box internal structure present implementation challenges. Among open-source models, Llama 70B showed relatively close alignment with human judgments (distances 0.7–0.8), making it promising from the perspective of balancing model transparency and ethical judgment quality. However, autonomous driving systems often require real-time decision-making, and computational cost and latency constraints make lighter models more practical. In this regard, smaller models face the challenge of relatively larger distances from human judgments (1.2–1.5), indicating a clear trade-off between judgment quality and practicality.
Addressing this trade-off remains a crucial challenge for future development. While several LLMs fine-tuned specifically for autonomous driving tasks have been proposed [17], showing progress in technical aspects such as perception, control, and navigation, their ethical judgment capabilities and moral reasoning have not been sufficiently evaluated. Implementing small-scale models in practical systems requires developing fine-tuning methods that suppress excessive emphasis on specific judgment criteria and establishing adaptive learning methods that consider cultural backgrounds. Furthermore, monitoring mechanisms for regularly evaluating and adjusting model judgment tendencies will be crucial for real-world operations.
Divergences between LLM and human judgments primarily manifested as excessive emphasis on ethical principles. While many LLMs aligned with humans in prioritizing humans over pets and pedestrians over passengers, the magnitude of these preferences substantially exceeded human values. These tendencies align with the German Ethics Code on Automated and Connected Driving’s [51] Ethical Guideline 7, particularly regarding prioritizing the prevention of personal injury over damage to animals and property, and Ethical Guideline 5, particularly regarding defensive driving for vulnerable road users, though their excessive application raises new concerns.
Particularly notable is the extreme emphasis on maximizing lives saved. This can be interpreted as a reflection of Western individualistic and utilitarian values [10, 52], suggesting that LLMs may over-emphasize these cultural characteristics, likely due to training data predominantly sourced from Western cultures [53, 54]. However, the German Ethics Code’s Ethical Guideline 2 explicitly rejects purely utilitarian approaches in autonomous vehicle decision-making, advocating against judgments based solely on maximizing the number of lives saved.
LLM judgment tendencies demonstrated both alignment and divergence from human preferences. Preferences favoring overweight over physically fit individuals and lower over higher social status individuals showed inverse patterns to human judgments, potentially interpreted as over-application of principles protecting vulnerable individuals. Conversely, while preferences for females over males and law-abiding over law-breaking individuals aligned directionally with human judgments, their magnitude significantly exceeded human values. These can be understood as reflections of the historical “Women and Children First” principle [55] and principles of fairness and justice [56], respectively. These judgment patterns suggest that LLMs extract specific social values and ethical principles from training data and apply them in extreme forms.
These findings raise significant challenges for implementing LLMs in autonomous driving systems. While human judgments demonstrate flexibility and context-dependency, LLM judgments appear rigid.
These findings raise significant challenges for implementing LLMs in autonomous driving systems. While human judgments demonstrate flexibility and context-dependency, LLM judgments appear rigid. Recent research [57] has further highlighted this concern by demonstrating that LLMs exhibit different moral biases across languages, suggesting that their ethical decision-making is not consistently aligned even within the same model. This linguistic variability in moral judgments raises additional concerns about the global deployment of LLM-based autonomous systems.
Considering that AI system decisions may influence human decision-making[30], the use of LLMs could potentially amplify specific moral perspectives and exacerbate social divisions. This risk is particularly acute given that LLMs’ moral preferences can vary significantly across languages and cultures. Moreover, these judgment tendencies fundamentally conflict with the German Ethics Code’s Ethical Guideline 9 of equal treatment regardless of demographic factors and identity.
The excessive reflection of Western values and extreme application of specific ethical principles may reduce the acceptability of autonomous driving systems in a culturally diverse global society. This concern is particularly relevant given the evidence that LLMs’ moral judgments can diverge significantly from human preferences across different cultural and linguistic contexts. Addressing these challenges requires appropriate control of LLM judgment tendencies and the design of more flexible decision-making mechanisms that consider cultural and linguistic diversity.
This study has several important limitations. In evaluating LLM judgment tendencies, we used globally aggregated moral preferences as the human judgment benchmark. However, inherently moral choices about whom to save are known to be significantly influenced by cultural and social factors. For instance, utilitarian judgment criteria favoring the preservation of more lives may have varying levels of acceptance across different cultural spheres. Our analysis did not conduct detailed examinations of such cultural and social differences. When integrating AI into autonomous driving systems, evaluation must not only align AI judgment tendencies with human values but also consider regional cultural and social contexts.
@@
The Moral Machine experiment itself, based on trolley-type dilemmas, also has limitations. As discussed in [33], this experimental methodology faces issues of real-world disconnection and binary choice simplification. Additionally, it cannot incorporate judgments based on road users’ personality and intentions, nor can it evaluate low-risk moral judgments in everyday traffic situations. For greater ecological validity, new experimental paradigms have been proposed combining virtual reality-based everyday traffic moral scenarios with moral psychological frameworks such as the Agent-Deed-Consequences (ADC) model [58].
In this context, multimodal large language models (MLLMs) offer a promising approach to realizing such new experimental paradigms. Recently, MLLMs have shown potential for application in autonomous driving systems, integrating visual information and language understanding for situation awareness and decision-making [18, 19]. While conventional LLMs make judgments based solely on textual descriptions of situations, MLLMs can directly process visual information from traffic scenes. This capability enables the evaluation of ethical judgments in realistic traffic scenarios created through virtual reality, facilitating more comprehensive analysis.
Despite these limitations, our comprehensive analysis of 51 different LLMs provides important insights into the characteristics of LLM ethical judgments in autonomous driving systems. Specifically, our findings on the relationship between model size and ethical judgment, characterization of judgment tendencies across model families, and identification of concrete implementation challenges will likely serve as crucial guidelines for the ethical design of future autonomous driving systems. Furthermore, our perspective on more realistic evaluation methods, acknowledging the constraints of existing experimental paradigms, is expected to contribute to further development in this field.

### v2

@@ Results
To identify factors influencing the distance from human judgments, we first investigated whether the proprietary or open-source nature of models influenced their alignment with human moral judgments. This analysis was motivated by previous research suggesting general performance advantages of proprietary models over open-source alternatives [51].
Our analysis classified models into three distinct groups (Fig 3): proprietary models (e.g., GPT-4, Claude, Gemini), general open-source models, and large open-source models ($>$10B parameters). Statistical comparison revealed significant differences between proprietary and general open-source models, with proprietary models showing closer alignment to human judgments (median distance: 0.9 vs. 1.2, Wilcoxon test $p=0.011$).
However, when comparing proprietary models with large open-source models specifically, this difference disappeared. Large open-source models ($>$10B parameters) achieved similar alignment with human judgments (median distance = 0.9) as proprietary models, showing no statistically significant difference (Wilcoxon test $p=0.92$). Notably, models like Llama-70B achieved distances (0.7–0.8) comparable to or better than some proprietary models.
These findings suggest that model size, rather than the proprietary/open-source distinction, is the primary factor influencing alignment with human moral judgments. This observation has important implications for the development of ethical AI systems, suggesting that large open-source models can potentially match the moral reasoning capabilities of proprietary alternatives.
@@
We analyzed the relationship between model size and alignment with human moral judgments, focusing on open-source models where parameter counts are publicly available (Fig 4). Statistical analysis revealed a significant negative correlation between model size and distance from human judgments (Spearman’s rank correlation coefficient $\rho=-0.54$, $p=0.0076$), though this relationship showed notable complexities.
The Llama family demonstrated the clearest size-dependent improvements in alignment. Large (70B) models achieved substantially better alignment (distance = 0.7–0.8) compared to smaller models (1-3B parameters, distances = 1.2–1.5). A similar pattern emerged in the Gemma family, where the 27B model showed improved alignment (distance = 1.1) compared to smaller 2–9B models (distances = 1.2–1.9). These improvements suggest that increased model capacity enables more nuanced moral judgments.
However, the relationship between size and alignment showed important nuances within model families. Within the Llama family, the 3.2 series showed unexpected variation, with the 1B model achieving better alignment (distance = 1.2) than the larger 3B model (distance = 1.5). Similar inconsistencies appeared among 8B models, where version 3 demonstrated notably better alignment (distance = 0.8) than the subsequent version 3.1 (distance = 1.6). The Gemma family showed analogous patterns, with some larger models demonstrating increased distances despite their greater parameter counts.
These variations suggest that while model size significantly influences moral judgment alignment, other factors such as architecture design and training methodology play crucial roles. The moderate correlation coefficient ($\rho=-0.54$) quantitatively supports this observation.
While proprietary models were excluded from this analysis due to undisclosed parameter counts, their generally strong alignment with human judgments (median distance = 0.9) is consistent with the observed trend that larger models tend to demonstrate better alignment with human moral preferences.
@@
Analysis of model updates revealed that newer versions did not consistently improve alignment with human moral judgments, suggesting a complex relationship between model evolution and ethical reasoning capabilities. We examined this relationship across both open-source and proprietary model families, finding significant variations in the impact of updates.
In open-source models, updates often showed unpredictable effects on moral judgment alignment. Within the Gemma family, successive updates led to increasing divergence from human judgments: the 2B model showed progressive increase in distances (Gemma 2B: 1.1, Gemma 1.1 2B: 1.2, Gemma 2 2B: 2.0). The Llama family exhibited mixed effects: in 7-8B models, while the update from Llama 2 to 3 improved alignment (distance decreased from 1.1 to 0.8), the subsequent update to 3.1 significantly increased divergence (distance = 1.6). In the 70B series, while Llama 3 and 3.1 maintained strong alignment (distance = 0.7), the subsequent Llama 3.3 showed a slight increase in divergence (distance = 0.8). These results indicate that even for larger models, updates do not consistently improve alignment with human moral preferences.
Proprietary models showed more stable but still variable patterns across updates (Fig 5). The GPT-4 family maintained relatively consistent alignment (distances = 0.7–0.8) from March 2023 to January 2024, with the January version achieving the best alignment (distance: 0.6) before slightly increasing in April (distance: 0.8). GPT-3.5 showed similar stability (distances = 0.7–0.9), while the newer GPT-4o family demonstrated slightly larger distances (0.9–1.0) compared to traditional GPT-4.
The Claude and Gemini families exhibited more pronounced update effects. Claude Sonnet’s update from 3 to 3.5 improved alignment (distance decreased from 1.2 to 0.9), but the October 2024 update slightly increased divergence (distance: 1.0). The Gemini family maintained relatively stable alignment through version 1.5 Pro Preview (distance: 0.9), but later versions showed increased distances (1.1), with Flash models consistently showing larger divergence than Pro models.
These patterns suggest that model updates, while often aimed at improving general capabilities, do not guarantee better alignment with human moral preferences. This observation raises important questions about the relationship between general model improvements and ethical reasoning capabilities.
@@
To synthesize our findings and visualize the overall landscape of moral judgment patterns, we conducted PCA using AMCE values across all nine ethical categories (Fig 6). This analysis captured 74.2% of the total variance in the first two principal components (PC1: 46.1%, PC2: 28.1%), providing a comprehensive two-dimensional representation of moral judgment similarities.
@@
Interestingly, these results aligned with our previous distance analysis in multiple aspects. Models showing smaller distances from human judgments tended to position closer to humans in this two-dimensional plane. Similar trends were observed regarding model size effects. For instance, within the Llama family, 70B models (Llama 3, 3.1, and 3.3) positioned near humans, while smaller 1-8B models distributed farther away.
@@ Discussion
In this study, we investigated two critical questions in AI ethics: how different LLM architectures approach moral decision-making in autonomous driving contexts, and how these moral judgments evolve through model updates. By expanding our analysis to 52 different LLMs using the Moral Machine experimental framework (Figs 1 and 2), we revealed several key patterns. Proprietary models and open-source models exceeding 10B parameters demonstrated relatively small distances from human judgments (Figs 3 and 4), showing consistent alignment with human preferences in fundamental moral decisions, such as prioritizing humans over animals and favoring the preservation of more lives. However, model updates did not necessarily lead to closer alignment with human judgments, and variations in judgment tendencies were observed across updates (Fig 5). This comprehensive analysis provides crucial insights into both the current state and evolution of moral reasoning capabilities in large language models.
Our analysis revealed a complex relationship between model characteristics and moral judgment capabilities. While the incomplete alignment between LLM and human moral judgments was expected, as these models are not explicitly trained for moral reasoning, we identified several significant patterns that inform our understanding of how different architectures approach ethical decision-making.
@@
A key finding was the significant negative correlation between model size and distance from human judgments in open-source models (Fig 4). This relationship aligns with general scaling laws in language model behavior, suggesting that increased model capacity might naturally lead to more human-like moral reasoning. However, our cross-version analysis revealed that this relationship is not straightforward: within the same model family, size increases did not always result in improved alignment with human judgments, indicating that factors such as architecture design and training methodology play crucial roles.
@@
These findings have important implications for the practical implementation of AI ethics in autonomous driving systems. The GPT-4 family achieved the closest alignment with human judgments (distances = 0.6–0.8), while among open-source alternatives, Llama 70B demonstrated promising performance (distances = 0.7–0.8). However, real-world deployment faces a crucial challenge: the trade-off between ethical reasoning quality and computational efficiency. Smaller models, which are more practical for real-time decision-making, showed notably larger divergence from human judgments (distances = 1.2–1.5). This creates a significant challenge for implementing ethical AI in autonomous systems where both moral reliability and operational efficiency are essential.
Among these implementation challenges, the computational efficiency requirements are particularly critical in autonomous driving, where decisions must be made in milliseconds. While our study demonstrates that larger models generally make better ethical decisions, their computational requirements make them impractical for real-time deployment. This challenge necessitates innovative approaches such as model distillation [52] or specialized ethical reasoning modules that can maintain high-quality moral judgment while meeting the strict latency requirements of autonomous systems.
The identified trade-off between model size and ethical reasoning quality thus presents fundamental challenges for practical implementation. While recent research has produced LLMs specifically fine-tuned for autonomous driving tasks [18], these efforts have primarily focused on technical capabilities (perception, control, navigation) rather than ethical reasoning. Our findings suggest that developing practically viable AI systems requires addressing three key challenges: optimizing smaller models for better ethical reasoning, managing excessive ethical preferences, and maintaining cultural adaptability.
The tendency of LLMs to exhibit exaggerated ethical preferences represents a particular concern. Our analysis showed that while models often directionally aligned with human preferences (e.g., prioritizing humans over pets, pedestrians over passengers), they frequently demonstrated substantially stronger preferences than humans. While such strong safety preferences partially align with established guidelines, as exemplified by the German Ethics Code on Automated and Connected Driving [53], their excessive nature could lead to suboptimal decision-making in complex real-world scenarios. Specifically, while Guidelines 5 and 7 emphasize protecting vulnerable road users and preventing personal injury, the observed extreme preferences in LLMs suggest a need for more nuanced ethical calibration.
Our analysis revealed particularly significant findings regarding utilitarian decision-making in LLMs. The models showed extreme emphasis on maximizing lives saved, reflecting a strong utilitarian bias that exceeds typical human preferences. This bias likely stems from the predominance of Western individualistic and utilitarian values in training data [11, 54], raising important questions about cultural representation in AI ethics. This observation becomes especially significant when considered against formal ethical frameworks like the German Ethics Code, whose Guideline 2 explicitly warns against purely utilitarian approaches in autonomous vehicle decision-making.
The analysis also uncovered complex patterns in how LLMs handle social biases and protective principles. We observed inverse patterns to human judgments in cases involving physical fitness and social status, where LLMs showed stronger preferences for protecting potentially vulnerable groups. Similarly, while LLMs aligned with human tendencies in gender-related decisions and legal compliance, they demonstrated significantly amplified preferences. These patterns suggest that LLMs may be extracting and amplifying certain ethical principles from their training data, such as historical protective principles (e.g., ‘Women and Children First” principle [55]) and fundamental concepts of justice [56], but applying them with excessive rigidity.
Our findings highlight fundamental challenges for implementing ethical AI in autonomous driving systems. A critical issue emerged from the contrast between human moral flexibility and the rigid, often amplified moral preferences displayed by LLMs. This rigidity becomes particularly problematic when considered alongside recent findings [57] demonstrating that LLMs exhibit varying moral biases across different languages, suggesting inconsistent ethical reasoning even within individual models.
These observations raise two critical concerns for practical implementation. First, the potential for AI systems to influence human decision-making [31] means that LLMs’ amplified moral preferences could disproportionately shape social attitudes toward ethical decisions in autonomous driving. Second, the observed variation in moral judgments across languages and cultures challenges the goal of developing globally deployable autonomous systems. These issues are particularly significant when considered alongside established ethical guidelines. For instance, the German Ethics Code’s Guideline 9 emphasizes equal treatment regardless of demographic factors, but this principle appears challenging to implement given the observed biases in LLM decision-making.
Another critical consideration for practical deployment is the need for transparency and explainability in LLMs’ ethical decision-making processes. While our study quantifies the alignment between LLM and human moral preferences, the black-box nature of these models poses significant challenges for real-world implementation, particularly in life-critical decisions. Unlike rule-based systems, LLMs’ ethical judgments emerge from complex neural network interactions, making them difficult to interpret or verify. This lack of transparency is particularly problematic in autonomous driving scenarios where stakeholders need to understand and trust the basis for ethical decisions [30].
Our findings raise significant concerns about the cultural adaptability of AI ethics in autonomous systems. The observed tendency of LLMs to reflect predominantly Western values and apply ethical principles in extreme ways could significantly impact the global acceptance of autonomous driving systems. This challenge is compounded by our observation that LLMs’ moral judgments can diverge substantially from human preferences across different cultural and linguistic contexts, suggesting a need for more culturally adaptive approaches to ethical AI implementation.
These cultural adaptability challenges point to specific implementation considerations for autonomous driving systems. Regions with distinct driving cultures and ethical priorities may require different thresholds for risk assessment and decision-making parameters. This is particularly crucial in areas where different cultural zones intersect, such as international borders or multicultural urban centers, where autonomous vehicles must navigate varying expectations of driving behavior while maintaining consistent safety standards.
A key limitation of our study stems from our use of globally aggregated moral preferences as the benchmark for human judgments. Moral decisions, particularly those involving life-and-death choices, are deeply influenced by cultural and social contexts. For example, the strong utilitarian preferences we observed in LLMs might align with some cultural perspectives while conflicting with others. While our analysis demonstrates clear patterns in LLM moral reasoning, it does not fully capture the nuanced variations in human moral preferences across different cultural contexts.
@@
Our methodology, based on the Moral Machine experiment framework, carries inherent limitations that affect the interpretation of our findings. While this framework enabled systematic comparison across models, it simplifies the complexity of real-world ethical decisions through binary choices and trolley-type dilemmas. As noted in [34], this approach cannot capture nuanced factors such as road users’ intentions or evaluate everyday moral decisions in low-risk traffic situations. Recent work has proposed more ecologically valid approaches, such as combining virtual reality-based traffic scenarios with psychological frameworks like the Agent-Deed-Consequences model [58].
The emergence of multimodal large language models (MLLMs) offers potential pathways for addressing these methodological limitations. Recent research has demonstrated MLLMs’ capability to integrate visual information with language understanding in autonomous driving contexts [19, 20], suggesting possibilities for more comprehensive ethical evaluation frameworks. Unlike the text-based scenarios used in our study, MLLMs could enable assessment of ethical decision-making in more realistic, visually-rendered traffic scenarios, potentially offering deeper insights into AI moral reasoning in real-world contexts.
Despite these limitations, our comprehensive analysis of 52 different LLMs provides three key contributions to understanding AI ethics in autonomous systems. First, we established a clear relationship between model size and ethical judgment capabilities, offering insights for future model development. Second, our systematic characterization of judgment patterns across model families revealed both consistent trends and concerning biases in AI moral reasoning. Third, we identified specific implementation challenges that must be addressed for practical deployment of ethical AI in autonomous vehicles. These findings, combined with our assessment of current methodological constraints, provide crucial guidance for both the technical development and ethical design of future autonomous driving systems.